what would de answer be for this system of equation, using substitution or elimination?
3x-2y=5
-6x+4y=7

Answer by stanbon(75887)

what would the answer be for this system of equation, using substitution or elimination?
3x-2y=5
-6x+4y=7
-----
Multiply the 1st equation by 2 to get:
6x-4y = 10
-6x+4y = 7
------------------
Add the equations to get:
0 = 17
---
That is a contradiction. There is
no solution for the system of equations.
If you graphed the two equations you would
see that the lines are parallel; they have
no point in common.
=========================
Cheers,
Stan H.